OpenStudy (jenniferjuice):

standard form of a linear equation? (1, 2) and (0, 3). m= -1

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ 0}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= -1 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and move everything to the left side}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jenniferjuice):

standard form?

OpenStudy (jenniferjuice):

@jdoe0001

OpenStudy (jdoe0001):

yeap http://www.mathsisfun.com/algebra/standard-form.html means you move everything to one side, either left or right and thus you'd equate it to 0 so you'd end up with something like 3x -25 = 0 or so

OpenStudy (jdoe0001):

well... hmm there's a y so it'll be more like 3x-4y-25=0 or so